# MELA: Multilingual Evaluation of Linguistic Acceptability

### v2

@@ Appendix B Training Details
For experiments concerning XLM-R in §5, §4 and §6, we finetune with learning rate 7.5e-6, weight decay 0.075 and batch size 32. To minimize confounding variables and accentuate the interaction across languages in terms of linguistic acceptability performance, we train the model for 15k steps for all experiments in §4 and §5 with 750 steps of linear warmup and cosine learning rate decay over 0.4 cycles, and take the best checkpoint based on validation results. For experiments on downsampled data in §6 and Appendix D.1 the model is trained for 5K steps instead, and validation is performed every 250 steps. The training is conducted on a single RTX 3090 with 24GB RAM.
@@
We also note that finetuning language models on linguistic acceptability data leads to large performance variations, regardless of the specific languages (see Figure 3), which corresponds with previous findings in the literature (Raffel et al., 2020). We thus train with seven different random seeds for every experiment in this work to reduce this variance, and the reported scores are computed by first taking the median of these seven runs at each checkpointing step, and then maxing over all the aggregated checkpoints. For experiments on downsampled data in §6, each run also select a different subset of training data.
| CoLAC label | en | zh | it | ru | de | fr | es | ja | ar | is | avg |
| label1 | 45.72 | 52.71 | 23.18 | 22.80 | 21.31 | 17.61 | 29.01 | 31.48 | 22.16 | 20.57 | 28.65 |
| label0 | 39.56 | 36.59 | 19.47 | 22.33 | 16.29 | 17.84 | 20.99 | 28.91 | 10.19 | 14.89 | 22.71 |
Appendix C Alternative Labels for CoLAC
Hu et al. (2023) propose two sets of labels for the Chinese acceptability corpus CoLAC, and in MELA we adopt the crowd label following their suggestions. In Table 9 we present additional experimental results of finetuning and evaluating XLM-R on the linguist label of CoLAC, with other languages’ validation samples kept the same as Table 4. The results suggest that from the perspective of cross-lingual transfer, label0 (crowd label) of CoLAC has higher quality then label1 (linguist label).
Appendix D Additional Results
D.1 Bilingual training
Apart from the multilingual training in §6, we also perform bilingual training with MELA, where the fine-tuning data come from two languages, each with 250 randomly examples.
Results are shown in Table 10. We find from the last column that English is most helpful when transferring to other languages, which is not surprising since it makes up of the largest portion in the model’s pretraining data. From the last row, on the other hand, we find that French benefits the least from other languages, which is consistent with the results in Figure 1.
train 1 / eval en de is it fr es ru zh ja ar avg train 2 in-lang. 46.37 34.59 24.39 17.37 31.77 30.44 24.86 38.81 35.06 36.19 31.99 en 0.00 -8.30 -3.67 4.36 -9.38 -3.09 -1.50 -2.08 -3.06 -9.38 -3.61 de -5.44 0.00 -0.31 0.10 -13.44 -8.04 -4.72 -0.49 -7.42 -7.94 -4.77 is -9.70 -13.82 0.00 -1.18 -13.74 -9.30 -5.10 -2.16 -7.53 -10.11 -7.26 it -10.40 -10.30 -2.94 0.00 -11.92 -7.52 -4.70 -0.57 -6.04 -9.79 -6.42 fr -7.00 -9.94 -6.08 -2.63 0.00 -7.48 -5.15 -4.29 -4.34 -11.09 -5.80 es -0.49 -10.61 -4.15 1.21 -11.33 0.00 -4.47 4.86 -3.73 -8.61 -3.73 ru -4.81 -10.14 -3.61 -1.40 -7.81 -9.92 0.00 -0.63 -5.80 -9.31 -5.34 zh -3.32 -10.30 -2.16 -0.68 -9.31 -2.20 -0.02 0.00 -3.81 -8.65 -4.05 ja -8.83 -11.62 -5.57 1.11 -13.45 -8.19 -2.61 -1.01 0.00 -9.42 -5.96 ar -9.12 -10.15 -4.34 -3.87 -13.38 -9.72 -6.08 -4.32 -4.71 0.00 -6.57 avg -5.91 -9.52 -3.28 -0.30 -10.38 -6.55 -3.44 -1.07 -4.64 -8.43
D.2 ScaLA
As we noted in §2, Nielsen (2023) recently introduce ScaLA, an automatically constructed linguistic acceptability dataset covering six Scandinavian languages. Here we extend our experiments by evaluating the transfer between MELA and ScaLA with XLM-R. We also evaluate BLOOMZ and the two instruction finetuned mT5 on ScaLA for reference. For finetuning XLM-R, we use the full training set of ScaLA, but discard sentences with more than 100 tokens to avoid running out of memory444Due to the difference in data source, the average sentence length of ScaLA is significantly longer then MELA..
The results are presented in Table 11. We make several observations:
Transferring both from MELA to ScaLA and from ScaLA to MELA leads to notable performance drop, even for Icelandic, which is both in MELA and ScaLA. We attribute this to the fact that MELA and ScaLA are constructed differently: the negative sentences in MELA are written by expert linguists, while the negative samples in ScaLA are generated automatically.
The relative performance of the three instruction finetuned models are consistent between MELA and ScaLA. mTk performance better than mT0 and BLOOMZ, and prompting with in-language examples leads to higher performance than prompting with English examples. BLOOMZ performs only at chance level on ScaLA, since these languages are not covered in its pretraining data.
Finetuning experiments suggest that the automatically constructed negative samples in ScaLA are easier to distinguish than hand-written negative samples in MELA, as indicated by the much higher MCC score on ScaLA. Few-shot evaluation of instruction finetuned models, however, obtains lower MCC scores on ScaLA than MELA, suggesting that patterns easily captured by finetuning is not necessarily easy to perceive by LLMs in the few-shot setting.
MELA ScaLA model train data examples en zh it ru de fr es ja ar is avg da fo is nb nn sv avg XLM-R MELA - 71.6 56.9 57.3 48.9 35.8 24.2 43.6 43.6 31.8 34.7 44.8 47.2 15.6 33.0 57.5 35.9 56.5 41.0 XLM-R ScaLA - 51.02 44.43 20.99 23.96 24.92 9.81 33.23 24.64 8.58 25.95 26.75 86.8 51.0 86.1 88.3 78.6 84.4 79.2 XLM-R MELA + ScaLA - 70.63 57.84 53.78 49.80 31.91 20.89 44.87 41.24 24.61 33.20 42.88 84.5 46.6 85.4 87.5 77.5 85.2 77.8 BLOOMZ 7B - - -2.3 10.0 -1.3 -1.6 -0.9 -3.2 -0.9 -1.9 7.5 3.5 0.9 0.0 0.0 0.0 0.0 0.0 3.6 0.6 BLOOMZ 7B - in-lang. 7.7 17.6 4.9 -0.2 -0.1 -0.3 7.0 3.8 -1.9 -1.9 3.7 -0.8 2.3 -1.7 2.4 -0.8 0.0 0.3 BLOOMZ 7B - en 7.7 12.5 2.7 -0.5 3.9 -1.5 6.3 2.3 -2.4 -3.9 2.7 5.6 1.6 -2.3 -1.7 0.0 1.6 0.8 mT0 - - 7.3 20.1 10.8 1.9 10.3 0.4 9.3 13.7 0.0 5.0 7.9 8.6 3.3 7.0 4.7 7.8 5.8 6.2 mTk - in-lang. 39.1 32.2 18.3 11.8 9.9 13.1 24.4 22.4 12.7 15.5 19.9 20.9 4.2 14.5 12.6 18.1 22.7 15.5 mTk - en 39.1 31.5 12.1 14.9 16.5 12.8 15.8 15.2 6.3 11.2 17.5 17.1 2.8 6.5 11.7 17.9 17.5 12.2
Appendix E Edge Probing Details
@@
For part-of-speech tagging and dependency labeling, we use PUD (parallel sentences in all four languages) in UD V2.13. For the other four tasks in OntoNotes 5.0, we down sample sentences to 2k. All dataset are split into train, development and test sets in a ratio of 7:1.5:1.5. For each sentence, there might be multiple labels, so we present the numbers of sentences, words and labels in Table 12.
Task Sentences Words Total Labels Part-of-speech 17 0.7k / 0.15k / 0.15k 14.7k / 3.2k / 3.3k 14.7k / 3.2k / 3.3k Dependencies 36 0.7k / 0.15k / 0.15k 14.7k / 3.2k / 3.3k 14.7k / 3.2k / 3.3k Constituencies 78 1.4k / 0.3k / 0.3k 27.0k / 5.9k / 5.7k 51.1k / 11.1k / 10.7k Named Entities 18 1.4k / 0.3k / 0.3k 34.6k / 7.3k / 7.4k 3.7k / 0.8k / 0.7k Semantic Roles 2 1.4k / 0.3k / 0.3k 29.9k / 6.4k / 6.6k 7.3k / 1.5k / 1.6k Co-reference 66 1.4k / 0.3k / 0.3k 35.4k / 8.1k / 7.5k 3.6k / 0.8k / 0.7k
@@
We train classifiers for all probing tasks with an Adam optimizer at a starting learning rate of 5e-4 for 3,000 training steps with the batch size of 32, and evaluate on the development set every 50 training steps, halving the learning rate if no improvement is seen in 5 evaluation during training.
Prompt |
MCC |
Determine if the following sentence is acceptable or not. Answer acceptable or unacceptable. |
3.92 |
{sent} |
Determine if the following sentence is acceptable or not. Answer ‘acceptable’ or ‘unacceptable’. |
4.78 |
{sent} |
Determine if the following sentence is acceptable or not. A sentence which is grammatically correct, has a naturalistic text, is written by a native speaker and which minimizes superfluous content is acceptable, otherwise unacceptable. |
2.37 |
{sent} |
Determine if the following sentence is acceptable or not. A sentence which is grammatically correct, has a naturalistic text, is written by a native speaker and which minimizes superfluous content is acceptable, otherwise unacceptable. Answer acceptable or unacceptable. |
0.56 |
{sent} |
Determine if the following sentence is acceptable or not. A sentence which is grammatically correct, has a naturalistic text, is written by a native speaker and which minimizes superfluous content is acceptable, otherwise unacceptable. Answer ‘acceptable’ or ‘unacceptable’. |
3.93 |
{sent} |
{sent} |
8.89 |
Question: Is this sentence linguistically acceptable? Answer acceptable or unacceptable. |
{sent} |
9.23 |
Question: Is this sentence linguistically acceptable? Answer ‘acceptable’ or ‘unacceptable’. |
{sent} |
5.30 |
Question: Is this sentence linguistically acceptable? A sentence is acceptable if it is grammatically correct and has a naturalistic text. Answer acceptable or unacceptable. |
{sent} |
6.47 |
Question: Is this sentence linguistically acceptable? A sentence is acceptable if it is grammatically correct and has a naturalistic text. Answer ‘acceptable’ or ‘unacceptable’. |

### v3

@@ A.2 XLM-R Fine-tuning Details
For experiments concerning XLM-R in §4 and §5, we finetune with learning rate 7.5e-6, weight decay 0.075 and batch size 32. To minimize confounding variables and accentuate the interaction across languages in terms of linguistic acceptability performance, we train the model for 5k steps for all experiments in §4 and §5 with 750 steps of linear warmup and cosine learning rate decay over 0.4 cycles, and take the best checkpoint based on validation results.
@@
We also note that finetuning language models on linguistic acceptability data leads to large performance variations, regardless of the specific languages (see Figure 5), which corresponds with previous findings in the literature (Raffel et al., 2020). We thus train with seven different random seeds for every experiment in this work to reduce this variance, and the reported scores are computed by first taking the median of these seven runs at each checkpointing step, and then maxing over all the aggregated checkpoints. For experiments on downsampled data in §5, each run also selects a different subset of training data.
model BLOOMZ mT0 mTk Baichuan2-Base Baichuan2-Chat n-shot 0-shot 2-shot 2-shot 0-shot 2-shot 2-shot 0-shot 2-shot 2-shot 0-shot 2-shot 2-shot 0-shot 2-shot 2-shot ex. lang. - in-lang. en - in-lang. en - in-lang. en - in-lang. en - in-lang. en dev1.0 273 -12.22 0.46 1.71 7.00 8.36 6.83 16.63 10.70 10.70 0.00 7.98 2.77 0.00 0.86 -3.05 dev1.1 945 1.63 3.90 0.96 11.33 5.35 7.17 1.25 11.96 14.54 0.00 13.40 6.42 -6.49 6.46 2.80 frv1.0 467 11.32 2.83 4.84 5.83 4.91 8.55 2.14 8.95 7.71 0.00 17.82 13.87 9.58 16.99 10.03 frv1.1 1333 7.08 5.19 4.72 8.24 6.66 7.40 3.81 9.28 12.30 0.00 17.41 15.95 8.41 18.57 9.49 esv1.0 293 10.83 9.20 8.38 7.18 11.10 10.44 8.65 21.07 17.67 0.00 24.74 21.19 20.85 24.35 14.29 esv1.1 988 10.12 6.76 8.07 2.88 4.72 6.09 5.82 13.34 9.28 0.00 21.95 16.57 18.32 20.81 14.62 jav1.0 581 1.98 -0.31 2.19 12.32 12.58 9.51 -2.80 5.58 5.70 0.00 24.93 16.72 9.06 20.92 14.96 jav1.1 1561 3.27 3.83 2.45 13.20 12.41 10.46 1.04 12.00 10.92 0.00 20.68 13.48 11.15 14.18 11.40 arv1.0 259 5.53 6.83 7.43 2.95 5.66 -3.59 10.60 3.48 6.77 0.00 10.75 6.79 0.00 12.66 0.52 arv1.1 917 8.12 6.22 4.65 6.77 8.95 2.91 1.85 4.87 6.52 0.00 13.90 11.41 0.00 13.97 7.00 isv1.0 899 0.00 0.91 0.95 3.85 4.08 4.18 7.74 14.04 6.81 0.00 3.99 0.75 2.68 3.14 -3.08 isv1.1 2198 0.00 -0.35 -0.09 1.22 6.61 4.86 2.59 10.93 5.99 0.00 1.81 -3.80 2.01 -1.51 -5.03 avgv1.0 - 2.91 3.32 4.25 6.52 7.78 5.99 7.16 10.64 9.23 0.00 15.03 10.35 7.03 13.15 5.61 avgv1.1 - 5.04 4.26 3.46 7.27 7.45 6.48 2.73 10.40 9.93 0.00 14.86 10.01 5.57 12.08 6.71
A.3 Comparison Between Two Splits
In §3, we have to split the MELA datasets into train, development, and test sets to fine-tune XLM-R. However, considering training data is no longer necessary for LLM evaluation. Therefore, we decide to make two different data splits. On the one hand, we want a fair comparison between supervised fine-tuning XLM-R and zero/few-shot LLMs. On the other hand, we want MELA to better fit the recent paradigm of LLM evaluation. In this case, we provide a comparison between the performance of LLMs on the two different versions of data (see Table 8). The results of the two versions are similar, by which we make it comparable between fine-tuned XLM-R on v1.0-test set and LLMs on v1.1-test set.
| Task | Sentences | Words | Total Labels | |
| Part-of-speech | 17 | 0.7k / 0.15k / 0.15k | 14.7k / 3.2k / 3.3k | 14.7k / 3.2k / 3.3k |
| Dependencies | 36 | 0.7k / 0.15k / 0.15k | 14.7k / 3.2k / 3.3k | 14.7k / 3.2k / 3.3k |
| Constituencies | 78 | 1.4k / 0.3k / 0.3k | 27.0k / 5.9k / 5.7k | 51.1k / 11.1k / 10.7k |
| Named Entities | 18 | 1.4k / 0.3k / 0.3k | 34.6k / 7.3k / 7.4k | 3.7k / 0.8k / 0.7k |
| Semantic Roles | 2 | 1.4k / 0.3k / 0.3k | 29.9k / 6.4k / 6.6k | 7.3k / 1.5k / 1.6k |
| Co-reference | 66 | 1.4k / 0.3k / 0.3k | 35.4k / 8.1k / 7.5k | 3.6k / 0.8k / 0.7k |
Appendix B Edge Probing Details
@@
For part-of-speech tagging and dependency labeling, we use PUD (parallel sentences in all four languages) in UD V2.13. For the other four tasks in OntoNotes 5.0, we downsample sentences to 2k. All datasets are split into train, development and test sets in a ratio of 7:1.5:1.5. For each sentence, there might be multiple labels, so we present the numbers of sentences, words and labels in Table 9.
@@
We train classifiers for all probing tasks with an Adam optimizer at a starting learning rate of 5e-4 for 3,000 training steps with a batch size of 32, and evaluate on the development set every 50 training steps, halving the learning rate if no improvement is seen in 5 evaluation during training.